OpenStudy (anonymous):

An equation is shown below: -2(4x - 1) - 7 = 5 Which statement shows a correct next step in solving the equation? The equation can become -2(4x - 1) = -2 by applying the distributive property. The equation can become -2(4x - 1) = 12 by applying the addition property of equality. The equation can become -2(4x - 1) = 12 by applying the commutative property of addition The equation can become -2(4x - 1) = -2 by applying the subtraction property of equality.

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@noseboy908 can you help

OpenStudy (noseboy908):

To solve this, think about what happened to the right side of the equation to move the "-7" over. What does one do to eliminate it from the left side?

OpenStudy (anonymous):

you would have to do the opposite

OpenStudy (noseboy908):

Yep. What is the opposite function of "-7"?

OpenStudy (anonymous):

7

OpenStudy (noseboy908):

You are correct, I think, but make sure to put the addition sign in front of it. It is assumed, but it's just good form to do so.

OpenStudy (anonymous):

ok srry

OpenStudy (noseboy908):

Not a problem, it simply helps with combining parts. Now, since you seem to know somewhat about simplifying these equations, how do you move the -7 to the right? What is done to the 5?

OpenStudy (anonymous):

?

OpenStudy (noseboy908):

Do what you did before: do the opposite of what's being done on a number to both sides. You're presently subtracting 7, so what's the opposite of that?

OpenStudy (anonymous):

i do not know

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

please help

OpenStudy (noseboy908):

You told me before what the opposite of -7 is. Just restate that.

OpenStudy (anonymous):

i still do not understand

OpenStudy (noseboy908):

Just restate what you said was the opposite function of "-7". You literally just need to scroll up.

OpenStudy (anonymous):

oh 7

OpenStudy (anonymous):

i mean +7 sorry

OpenStudy (noseboy908):

Yes. So knowing that, add 7 to both sides. What is the equation once you do so?

OpenStudy (anonymous):

i do not know

OpenStudy (noseboy908):

Just try to do as I instruct: Add 7 to both sides of "-2(4x - 1) - 7 = 5". On the left, you're adding 7 to -7. On the right, you're adding 7 to 5.

OpenStudy (anonymous):

-2(4x-1) - 7 = 5 + 7

OpenStudy (noseboy908):

Where's the 7 you add to the left side?

OpenStudy (anonymous):

i do not know where it is supposed to be

OpenStudy (noseboy908):

You just have to add it to the -7 already there. It cancels out.

OpenStudy (anonymous):

so show me where you put it

OpenStudy (noseboy908):



OpenStudy (anonymous):

ok

OpenStudy (noseboy908):

Now simplify the right side. Add the two values.

OpenStudy (anonymous):

is it 12

OpenStudy (noseboy908):

Correct, which leaves only two possible answers. What's left is to decide if we used the communicative property or the addition property. So ask yourself this: Did we simply show how when adding two values together, it does not matter what order they're in, or did we show how to add in an inequality?

OpenStudy (anonymous):

we simplified how when adding two values together

OpenStudy (noseboy908):

You might have it, I just don't understand how you worded it. Which of the two possibilities did we do? Try to word it as I have, for clarity sake.

OpenStudy (anonymous):

communicative property i think right

OpenStudy (noseboy908):

Did we show how if A + B can be rewritten as B+A, for example? That is the communicative property.

OpenStudy (anonymous):

ok?

OpenStudy (noseboy908):

I ask again, did we show how the order of a addition problem does not matter? Yes or no?

OpenStudy (anonymous):

no

OpenStudy (noseboy908):

Correct. And that is the definition of the communicative property. Thus, what is our answer?

OpenStudy (anonymous):

no

OpenStudy (noseboy908):

no what????

OpenStudy (anonymous):

oh our answer communicative property

OpenStudy (noseboy908):

We just eliminated that as a solution.

OpenStudy (anonymous):

oh so it is addition property

OpenStudy (noseboy908):

Yes, but I fear that you guessed. Can you explain why it's the addition property?

OpenStudy (anonymous):

because we showed the order of the addition problem does not matter

OpenStudy (anonymous):

i mean does sorry

OpenStudy (noseboy908):

No, that's the reason that it was not communicative. Why is it addition?

OpenStudy (anonymous):

i don't know then to be honest i just felt that it was

OpenStudy (noseboy908):

Think about it, what did we do to simplify the equation in the beginning? What function?

OpenStudy (anonymous):

we used addition

OpenStudy (noseboy908):

Yes. That, in short, is why we used the addition property. We'd of used the comm. property if we showed how the parts of an addition problem can be moved w/o changing the answer, but we didn't.

OpenStudy (anonymous):

ok thank you may i ask a few more questions please

OpenStudy (noseboy908):

Ask.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 2x - one over four (12x + 4) = 3 Step 1: 2x + 3x + 1 = 3 Step 2: 5x + 1 = 3 Step 3: 5x + 1 - 1 = 3 - 1 Step 4: 5x = 2 In which step was the error made and why? Step 1, because incorrect signs have been used Step 2, because the terms have been added incorrectly Step 3, because a number has been subtracted from both sides of the equation No error was made

OpenStudy (anonymous):

are you still there

OpenStudy (noseboy908):

Yep

OpenStudy (noseboy908):

\[2x-(1/4)(12x+4)=3\]

OpenStudy (noseboy908):

Is that it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you have to answer the question that it says

OpenStudy (noseboy908):

I am aware of that. I'm asking if that is the original equation.

OpenStudy (anonymous):

yes

OpenStudy (noseboy908):

Then start by simplifying. What do you get after you distribute? Try to not look at their steps.

OpenStudy (anonymous):

do parentheses first

OpenStudy (noseboy908):

Yes, but what is the result?

OpenStudy (anonymous):

i don't know i am not very good at these kind of equations

OpenStudy (noseboy908):

When you distribute, you have to, in this equation, apply the multiplication factor (1/4 here) to all parts within the parenthesis. Example to follow.

OpenStudy (anonymous):

ok

OpenStudy (noseboy908):

\[1/2(10x-8) distributes \to5x-4\]

OpenStudy (anonymous):

ok

OpenStudy (noseboy908):

So try to distribute 1/4 to 12x+4. What is the result?

OpenStudy (anonymous):

i need help

OpenStudy (noseboy908):

Distribute by multiplying each part of "12x+4" by 1/4. AKA, divide each part by 4.

OpenStudy (anonymous):

ok so it would be 12x+1

OpenStudy (noseboy908):

You got the second part right, but make sure to divide the 12 too.

OpenStudy (anonymous):

3x+1

OpenStudy (anonymous):

right

OpenStudy (noseboy908):

Completely. Now please restate the equation I wrote out, with 3x+1 in place of 1/4(...)

OpenStudy (anonymous):

ok 2x-1/4(3x+1)=3

OpenStudy (noseboy908):

No, we already distributed teh 1/4, so it's gone now.

OpenStudy (anonymous):

ok so take that out and that would be the equation right there right?

OpenStudy (noseboy908):

Yes. Please restate it as so.

OpenStudy (anonymous):

i am very sorry i am in a hurry may you please tell me the answers flat out i am not trying to take advantage but please i am in a rush to finish this, because i have to do something really important to me

OpenStudy (noseboy908):

I'll never just given answers, but I can give you the steps for finishing this one: Rewrite the equation as we've gotten it thus far, after distributing. Look for errors. If there is none, then keep simplifying until you either find one or decide there is none.

OpenStudy (anonymous):

its fine i will just have to wait sadly lets just keep on going

OpenStudy (noseboy908):

Then go ahead and restate the equation after you distribute.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x-(3x+1)=3